OpenStudy (anonymous):

consider the function y(t)=3os(t)-4 find the period and the amplitude

OpenStudy (anonymous):

cos

OpenStudy (anonymous):

A Cos[pt]-b A= amplitude period= 2pi/p

OpenStudy (anonymous):

as per the problem, 2*pi*f = 1 f=1/2*pi Period = T = 1/f = 2pi Amplitude : maximum = -1; minimum = -7

OpenStudy (anonymous):

what?? these are my option for period a)2pi b)4pi c)2/pi d)2 e)pi and for amplitude a)7pi b)2pi c)3pi d)3 e)-3 f)6

OpenStudy (anonymous):

Period is 2pi. Since this has a cos term amplitude varies periodically. However if the question was 3cos(t-4) then the amplitude is 3.